a homeowner estimates that it will take her 9 days to roof her house. A professional roofer estimates that he could roof the house in 4 days. how long will it take if the homeowner helps the roofer?
Found 2 solutions by HyperBrain, Alan3354:
Answer by HyperBrain(694)

Let x=no. of days it will take the two to finish the job.
...
After x days, the homeowner has done x/9 of the job, while the roofer has done x/4 of the job. Furthermore,
(x/4)+(x/9)=1
multiply both sides by 36.
9x+4x=36
13x=36
x=36/13=2 10/13 days
Power up,
HyperBrain!
visit my blog here

Answer by Alan3354(69443)

a homeowner estimates that it will take her 9 days to roof her house. A professional roofer estimates that he could roof the house in 4 days. how long will it take if the homeowner helps the roofer?
-------------------
A shortcut method for 2 elements is
product/sum = 4*9/(4+9)
=36/13